# A Complete and Bi-Continuous Invariant of Protein Backbones under Rigid Motion

### v2

@@ 4 Lipschitz bi-continuity of the invariant $\mathrm{BRI}$
The triangle inequality for the Euclidean distance implies that $|{\bf\it u}|\leq|{\bf\it u}-{\bf\it v}|+|{\bf\it v}|$, so $|{\bf\it u}|-|{\bf\it v}|\leq|{\bf\it u}-{\bf\it v}|$. Swapping the vectors, we get $|{\bf\it v}|-|{\bf\it u}|\leq|{\bf\it u}-{\bf\it v}|$. Combining the inequalities ±$\pm(|{\bf\it u}|-|{\bf\it v}|)\leq|{\bf\it u}-{\bf\it v}|$, we conclude that $|\,|{\bf\it u}|-|{\bf\it v}|\,|\leq|{\bf\it u}-{\bf\it v}|$ as required. ∎
@@
$\leq|({\bf\it u}^{\prime}-{\bf\it u})\cdot{\bf\it v}^{\prime}|+|{\bf\it u}\cdot({\bf\it v}^{\prime}-{\bf\it v})|\leq\leq|{\bf\it u}^{\prime}-{\bf\it u}|\cdot|{\bf\it v}^{\prime}|+|{\bf\it u}|\cdot|{\bf\it v}^{\prime}-{\bf\it v}|=$
@@
$|y(N^{\prime}_{i})-y(N_{i})=|\overrightarrow{C^{\prime}_{i-1}N^{\prime}_{i}}\cdot{\bf\it v}^{\prime}_{i}-\overrightarrow{C_{i-1}N_{i}}\cdot{\bf\it v}_{i}|\leq$
@@
Because the metric $L_{\infty}$ between invariants $\mathrm{BRI}$ ($m\times 9$ matrices) can be computed in linear time $O(m)$, Theorem 4.1 also completes condition (1.2f) in Problem 1.2. Theorem 4.8 will prove condition in 1.2(d).
@@
For the resulting motion $f$, we will prove that the atoms of $S$ are $\mu\delta$-close to the corresponding atoms of the image of $Q$, which we still denote by $N^{\prime}_{i},A^{\prime}_{i},C^{\prime}_{i}$ for simplicity. Because the atom $N^{\prime}_{1}$ is in the $x$-axis through $\overrightarrow{A_{1}N_{1}}$, the first basis vectors of length 1 coincide (${\bf\it u}^{\prime}_{1}={\bf\it u}_{1}$) and hence also uniquely define the other basis vectors (${\bf\it v}^{\prime}_{1}={\bf\it v}_{1}$, ${\bf\it w}^{\prime}_{1}={\bf\it w}_{1}$).
@@
Because the first bases coincide, we consider the second residue:
@@
[…]
@@ 5 Average invariant, diagrams, barcodes
Figure 3: Row 1: the Lipschitz continuity of $\mathrm{BRI}$ from Theorem 4.1 is illustrated on the left by perturbing hemoglobins in Example 4.7, whose main chains A of 141 residues are shown in the middle (oxygenated 1hho in green, standard 2hhb in cyan) and eight $\alpha$-helices found by [27] on the right. Row 2: the Backbone Invariant Diagram ($\mathrm{BID}$) of the hemoglobins 1hho vs 2hhb in the PDB, see Definition 5.1. Row 3: the Backbone Invariant Barcode ($\mathrm{BIB}$), see Example 5.2.
@@
More importantly, a rigidly repeated pattern such as $\alpha$-helix or $\beta$-strand has constant invariants over several residue indices, which are easily detectable in $\mathrm{BID}$ and visible in $\mathrm{BIB}$ as intervals of uniform color. The PDB uses the baseline algorithm DSSP (Define Secondary Structure of Proteins) [27], which depends on several manual parameters and sometimes outputs $\alpha$-helices of only two residues. For instance, the PDB files 1hho and 2hhb in Fig. 3 (right) include HELX_P4 consisting of only residues 50 and 51, and HELX_P5 of length 20 over residue indices $i=52,\dots,71$. Fig. 3 shows that a ‘constant’ interval of little noise appears only for $i=54,\dots,70$. Hence new invariants allow a more objective detection of secondary structures, which will be explored in future work.
While the complete $\mathrm{BRI}(S)$ can be used to compare backbones of the same length, the average invariant $\mathrm{Brain}(S)\in\mathbb{R}^{9}$ and the standard deviation invariant can help to visualize all backbones of different lengths on the same heatmap. In each image of Fig. 4, any protein backbone is represented by a single point $(x,y)$ whose coordinates are the two simplest statistics (average and standard deviation) of a fixed invariant across all residues in a fixed chain. The top images in Fig. 4 show that the deviations of all three invariants from Definition 3.1 can be as large as $0.25\textup{\AA}$. So the shapes of residue triangles can substantially vary even for a fixed backbone, while AlphaFold2 [26] assumes that they all have identical shapes.
@@ 6 Duplicates with identical coordinates
In a similar case [58], when five pairs of unexpected duplicates were found in the Cambridge Structural Database (CSD). Their integrity office agreed that a single atomic replacement should perturb geometry at least slightly, so all coordinates cannot remain the same. Five journals started investigations into the data integrity of the relevant publications [13].
@@
The recently published method [18] didn’t report any duplicates. Right now anyone can download the PDB files from Table 2 and see all coincidences of $x,y,z$ coordinates with their own eyes without any computations. Here are the links to the identical files in the first row of Table 2, where the 4-letter PDB id can be replaced with any other id: https://files.rcsb.org/download/1A0T.cif and
@@
One potential explanation of identical coordinates is the molecular replacement method [50], which uses an existing protein structure, often a previous PDB deposit or part thereof, to solve a new structure. If the newly calculated electron density map does not allow for further refinement then the coordinates may (reasonably) remain unchanged.
The same coincidences can happen with lower-quality cryo-EM maps in which an existing PDB structure may be placed but where the resolution may not allow for further refinement of atomic coordinates [44, 21].
We have checked that the found duplicate backbones also have identical distance matrices on $3m$ ordered atoms, which were slower to compute in time $O(m^{2})$ over two days on a similar machine. The widely used DALI server [19] also confirmed the found duplicates by the traditional Root Mean Square Deviation (RMSD) through optimal alignment. The DALI took about 30 min on average to find a short list of nearest neighbors of one chain in the whole PDB. Extrapolating this time to all pairwise comparisons for 707K+ cleaned chains yields 40+ years, slower by orders of magnitude than 6 hours needed for all comparisons of the complete invariants $\mathrm{BRI}$ on the same desktop computer. Our implementation of $\mathrm{RMSD}$ for Fig. 5 has the median time of 534 microseconds per pair of chains (of the same length), about 10 times slower than $L_{\infty}$ on $\mathrm{BRI}$s.
The FoldSeek algorithm [55] is claimed to be 4000 times faster than RMSD by Dali due to optimal alignments of 3-residue subchains instead of full original chains, which takes 3.65 days by the estimates above, still an order of magnitude slower than $L_{\infty}$ on $\mathrm{BRI}$s. But any similarity distance needs a proof of all metric axioms for trustworthy clustering [49].
@@ 7 Discussion of the PDB and data integrity
The Lipschitz bi-continuity is essential because small perturbations of input should not drastically change the output and vice versa. Weaker versions of Problem 1.2 were solved for low-dimensional lattices [10, 33, 32], periodic point sets [16, 2, 4], sequences [5, 6, 31], crystals [56, 57], and finite clouds of unordered points [34, 59] within Geometric Data Science [3, 7].
@@
While traditional approaches explored finite datasets within infinite spaces in a ‘horizontal’ way, solutions to Problem 1.2 and its analogs for other data [60, 61] provide ‘vertical’ breakthroughs by building geographic-style maps of continuous data spaces as viewed from a satellite [9].
Fig. 2 and 4 can be zoomed at any spot and mapped by using further invariants. Such navigation maps with invertible coordinates enable inverse design while any dimensionality reduction to a latent space was proved [35] to be discontinuous (making close points distant) or collapsing an unbounded region to a point (losing an infinite amount of data).
@@
The supplementary data (available by request) include the Python code and a table of exact duplicates whose corresponding coordinates coincide in all decimal places and hence might need further refinement. Improving the PDB validation is needed to avoid unjustified predictions and claims of ‘solutions’ based on skewed data [41]. Another analysis [63] raised further concerns by revealing large numbers of waterless structures in the PDB.
@@ A Appendix: updates on PDB duplicates
Because all $x,y,z$ coordinates in the PDB are given with three decimal places relative to $1\textup{\AA}$, a distance of less than $0.01\textup{\AA}$ is considered negligible.

### v3

@@ 4 Lipschitz bi-continuity of the invariant $\mathrm{BRI}$
The triangle inequality for the Euclidean distance implies that $|{\bf\it u}|\leq|{\bf\it u}-{\bf\it v}|+|{\bf\it v}|$, so $|{\bf\it u}|-|{\bf\it v}|\leq|{\bf\it u}-{\bf\it v}|$. Swapping the vectors, we get $|{\bf\it v}|-|{\bf\it u}|\leq|{\bf\it u}-{\bf\it v}|$. Combining the inequalities $\pm(|{\bf\it u}|-|{\bf\it v}|)\leq|{\bf\it u}-{\bf\it v}|$, we conclude that $|\,|{\bf\it u}|-|{\bf\it v}|\,|\leq|{\bf\it u}-{\bf\it v}|$ as required. ∎
@@
$\leq|({\bf\it u}^{\prime}-{\bf\it u})\cdot{\bf\it v}^{\prime}|+|{\bf\it u}\cdot({\bf\it v}^{\prime}-{\bf\it v})|\leq|{\bf\it u}^{\prime}-{\bf\it u}|\cdot|{\bf\it v}^{\prime}|+|{\bf\it u}|\cdot|{\bf\it v}^{\prime}-{\bf\it v}|=$
@@
$|y(N^{\prime}_{i})-y(N_{i})|=|\overrightarrow{C^{\prime}_{i-1}N^{\prime}_{i}}\cdot{\bf\it v}^{\prime}_{i}-\overrightarrow{C_{i-1}N_{i}}\cdot{\bf\it v}_{i}|\leq$
@@
Since the metric $L_{\infty}$ between invariants $\mathrm{BRI}$ ($m\times 9$ matrices) can be computed in linear time $O(m)$, Theorem 4.1 also completes condition (1.2f) in Problem 1.2. Theorem 4.8 will prove condition in 1.2(d).
@@
For the resulting motion $f$, we will prove that the atoms of $S$ are $\mu\delta$-close to the corresponding atoms of the image of $Q$, which we still denote by $N^{\prime}_{i},A^{\prime}_{i},C^{\prime}_{i}$ for simplicity. Since the atom $N^{\prime}_{1}$ is in the $x$-axis through $\overrightarrow{A_{1}N_{1}}$, the first basis vectors of length 1 coincide (${\bf\it u}^{\prime}_{1}={\bf\it u}_{1}$) and hence also uniquely define the other basis vectors (${\bf\it v}^{\prime}_{1}={\bf\it v}_{1}$, ${\bf\it w}^{\prime}_{1}={\bf\it w}_{1}$).
@@
Since the first bases coincide, we consider the second residue:
@@
 $\text{ where }L=\max\{L_{C,N},L_{N,A},L_{A,C}\},\quad K=\dfrac{1}{l_{N,A}}+\dfrac{2}{h}\Big(1+2\dfrac{L_{A,C}}{l_{N,A}}\Big).$
@@ 5 Average invariant, diagrams, barcodes
Figure 3: Row 1: the Lipschitz continuity of $\mathrm{BRI}$ from Theorem 4.1 is illustrated on the left by perturbing hemoglobins in Example 4.7, whose main chains A of 141 residues are shown in the middle (oxygenated 1hho in green, standard 2hhb in cyan) and eight $\alpha$-helices found by [kabsch1983dictionary] on the right. Row 2: the Backbone Invariant Diagram ($\mathrm{BID}$) of the hemoglobins 1hho vs 2hhb in the PDB, see Definition 5.1. Row 3: the Backbone Invariant Barcode ($\mathrm{BIB}$), see Example 5.2.
@@
More importantly, a rigidly repeated pattern such as $\alpha$-helix or $\beta$-strand has constant invariants over several residue indices, which are easily detectable in $\mathrm{BID}$ and visible in $\mathrm{BIB}$ as intervals of uniform color. The PDB uses the baseline algorithm DSSP (Define Secondary Structure of Proteins) [kabsch1983dictionary], which depends on several manual parameters and sometimes outputs $\alpha$-helices of only two residues. For instance, the PDB files 1hho and 2hhb in Fig. 3 (right) include HELX_P4 consisting of only residues 50 and 51, and HELX_P5 of length 20 over residue indices $i=52,\dots,71$. Fig. 3 shows that a ‘constant’ interval of little noise appears only for $i=54,\dots,70$. Hence new invariants allow a more objective detection of secondary structures, which will be explored in future work.
While the complete $\mathrm{BRI}(S)$ can be used to compare backbones of the same length, the average invariant $\mathrm{Brain}(S)\in\mathbb{R}^{9}$ and the standard deviation invariant can help to visualize all backbones of different lengths on the same heatmap. In each image of Fig. 4, any protein backbone is represented by a single point $(x,y)$ whose coordinates are the two simplest statistics (average and standard deviation) of a fixed invariant across all residues in a fixed chain. The top images in Fig. 4 show that the deviations of all three invariants from Definition 3.1 can be as large as $0.25\textup{\AA }$. So the shapes of residue triangles can substantially vary even for a fixed backbone, while AlphaFold2 [jumper2021highly] assumes that they all have identical shapes.
@@ 6 Duplicates with identical coordinates
In a similar case [widdowson2022resolving], when five pairs of unexpected duplicates were found in the Cambridge Structural Database (CSD). Their integrity office agreed that a single atomic replacement should perturb geometry at least slightly, so all coordinates cannot remain the same. Five journals started investigations into the data integrity of the relevant publications [chawla2024crystallography].
@@
The recently published method [guzenko2020real] didn’t report any duplicates. Right now anyone can download the PDB files from Table 2 and see all coincidences of $x,y,z$ coordinates with their own eyes without any computations. Here are the links to the identical files in the first row of Table 2, where the 4-letter PDB id can be replaced with any other id: https://files.rcsb.org/download/1A0T.cif and
@@
One potential explanation of identical coordinates is the molecular replacement method [rossmann1990molecular], which uses an existing protein structure, often a previous PDB deposit or part thereof, to solve a new structure. If the newly calculated electron density map does not allow for further refinement then the coordinates may (reasonably) remain unchanged.
The same coincidences can happen with lower-quality cryo-EM maps in which an existing PDB structure may be placed but where the resolution may not allow for further refinement of atomic coordinates [murshudov2011refmac5, hekkelman2024pdbredo].
We have checked that the found duplicate backbones also have identical distance matrices on $3m$ ordered atoms, which were slower to compute in time $O(m^{2})$ over two days on a similar machine. The widely used DALI server [holm2024dali] also confirmed the found duplicates by the traditional Root Mean Square Deviation (RMSD) through optimal alignment. The DALI took about 30 min on average to find a short list of nearest neighbors of one chain in the whole PDB. Extrapolating this time to all pairwise comparisons for 707K+ cleaned chains yields 40+ years, slower by orders of magnitude than 6 hours needed for all comparisons of the complete invariants $\mathrm{BRI}$ on the same desktop computer. Our implementation of $\mathrm{RMSD}$ for Fig. 5 has the median time of 534 microseconds per pair of chains (of the same length), about 10 times slower than $L_{\infty}$ on $\mathrm{BRI}$s.
The FoldSeek algorithm [van2024fast] is claimed to be 4000 times faster than RMSD by Dali due to optimal alignments of 3-residue subchains instead of full original chains, which takes 3.65 days by the estimates above, still an order of magnitude slower than $L_{\infty}$ on $\mathrm{BRI}$s. But any similarity distance needs a proof of all metric axioms for trustworthy clustering [rass2024metricizing].
@@ 7 Discussion of the PDB and data integrity
The Lipschitz bi-continuity is essential because small perturbations of input should not drastically change the output and vice versa. Weaker versions of Problem 1.2 were solved for low-dimensional lattices [bright2023continuous, kurlin2024mathematics], periodic point sets [edelsbrunner2021density, anosova2021isometry, anosova2026recognition], sequences [anosova2022density, anosova2023density, kurlin2025complete], crystals [widdowson2022average, mcmanus2025computing, widdowson2025higher, widdowson2026pointwise], and finite clouds of unordered points [kurlin2024polynomial, widdowson2023recognizing] in Geometric Data Science [anosova2021introduction, anosova2025geometric, anosova2024importance].
@@
While traditional approaches explored finite datasets within infinite spaces in a ‘horizontal’ way, solutions to Problem 1.2 and its analogs for other data [kurlin2023strength, kurlin2023simplexwise, anosova2026seeing] provide ‘vertical’ breakthroughs by building geographic-style maps of continuous data spaces as viewed from a satellite [bright2023geographic, widdowson2024continuous, widdowson2025geographic].
Fig. 2 and 4 can be zoomed at any spot and mapped by using further invariants. Such navigation maps with invertible coordinates enable inverse design while any dimensionality reduction to a latent space was proved [landweber2016fiber] to be discontinuous (making close points distant) or collapsing an unbounded region to a point (losing an infinite amount of data).
@@
The supplementary data (available by request) include the Python code and a table of exact duplicates whose corresponding coordinates coincide in all decimal places and hence might need further refinement. The follow-up work [wlodawer2025duplicate] discusses specific duplicates in detail. Improving the PDB validation is needed to avoid unjustified predictions and claims of ‘solutions’ based on skewed data [mcdonnell2024structure]. Another analysis [wlodawer2024waterless] raised further concerns by revealing large numbers of waterless structures in the PDB.
@@ A Appendix: updates on PDB duplicates
Since all $x,y,z$ coordinates in the PDB are given with three decimal places relative to $1\textup{\AA }$, a distance of less than $0.01\textup{\AA }$ is considered negligible.